# Elucidating the Design Space of Multimodal Protein Language Models

### v2

@@ Abstract
Multimodal protein language models (PLMs) integrate sequence and token-based structural information, serving as a powerful foundation for protein modeling, generation, and design. However, the reliance on tokenizing 3D structures into discrete tokens causes substantial loss of fidelity about fine-grained structural details and correlations. In this paper, we systematically elucidate the design space of multimodal PLMs to overcome their limitations. We identify tokenization loss and inaccurate structure token predictions by the PLMs as major bottlenecks. To address these, our proposed design space covers improved generative modeling, structure-aware architectures and representation learning, and data exploration. Our advancements approach finer-grained supervision, demonstrating that token-based multimodal PLMs can achieve robust structural modeling. The effective design methods dramatically improve the structure generation diversity, and notably, folding abilities of our 650M model by reducing the RMSD from 5.52 to 2.36 on PDB testset, even outperforming 3B baselines and on par with the specialized folding models.
@@ 2.1 Multimodal generative learning for proteins with language models and structure tokenization
Diffusion protein language model (DPLM, Wang et al., 2024b), in particular, shows excelling performance in both generation and representation learning of protein sequences, and even structures thanks to its recent multimodal extension DPLM-2 (Wang et al., 2024a). The family of DPLMs grounded in absorbing discrete diffusion framework (Austin et al., 2021; Zheng et al., 2023), which is characterized by a forward and backward Markov process. Let $\texttt{Cat}(\mathbf{x};\mathbf{p})$ be a categorical distribution on protein sequence $\mathbf{y}$ parameterized by a vector $\mathbf{p}$ on $(|\mathcal{V}|-1)$-dimensional probability simplex. The forward process of discrete diffusion defines a Markov process governed by the transition kernel $q(\mathbf{x}^{(t)}|\mathbf{x}^{(t-1)})=\texttt{Cat}\big{(}\mathbf{x}^{(t)};\beta_{t}\mathbf{x}^{(t-1)}+(1-\beta_{t})\mathbf{q}_{\text{noise}}\big{)}$ that gradually perturb the data $\mathbf{x}\sim q(\mathbf{x})$ into a stationary distribution $\mathbf{x}^{(T)}\sim\mathbf{q}_{\text{noise}}$. The learned backward process $p_{\mathbf{\theta}}(\mathbf{x}^{(t-1)}|\mathbf{x}^{(t)})$ reversely denoises the $\mathbf{x}^{(T)}$ towards the data distribution $\mathbf{x}$, which is typically optimized by the variational bound of the log-likelihood (Ho et al., 2020). The learning objective of absorbing diffusion can be simplified into weighted cross-entropiess, resembling masked language modeling at arbitrary noise levels:
@@ 7 Conclusions
Acknowledgements
We would like to especially thank Dr. Hang Li for insightful discussions on the project and feedback on the manuscript that help shape this study. We thank Yi Zhou, Jing Yuan and Yilai Li for their valuable comments.
@@ Appendix B Discussions and Limitations
  • Third, our analysis is conducted on relatively small protein language models (up to 3B), and the scalability of these design choices remains uncertain. We chose this setup to prioritize the analysis of our design space. Further study into how these designs scale with larger models could further enhance multimodal protein language modeling, especially given the proven benefits of scaling in DPLM2.
@@
Ultimately, the future of multimodal protein foundation models should enable direct atom-level data-space structure modeling; however, this remains an unsolved challenge. Token-based multimodal protein LMs provide a scalable, practical, and effective approach to generative protein modeling. Advancing this approach while addressing its limitations—such as structural information loss, lack of physical constraints, and limited multimodal generalization—will be key to unlocking their full potential. Advancing this approach while addressing its limitations is the key to unlocking their full potential and provides crucial insights that help us move closer to achieving atomic-resolution protein modeling in a unified multimodal framework.
@@ B.1 Discussion on Alternative Approaches
  • Alternatives: Training DPLM2 on continuous tokens. However, it’s unclear whether this approach would work for joint modeling of structures and sequence, which is inherently discrete.
@@ C.2 Bit-level Supervision
In the training stage, we follow Wang et al. (2024a) and utilize the pre-trained DPLM (Wang et al., 2024b) as the parameter initialization to inherit the evolutionary representations learned from the massive protein sequence data, which implicitly captures the structural information that benefits structure modeling (Lin et al., 2022). In addition to the parameters inherited from DPLM, we randomly initialize the structure input-output linear layers. The bit-level DPLM-2 is grounded in the absorbing discrete diffusion framework (Austin et al., 2021; Zheng et al., 2023). To adapt to the absorbing discrete diffusion framework, we also introduce a learnable absorbing embedding for the absorbing state of structure. We employ 2,000 warmup steps until reaching the maximum learning rate $1\times 10^{-4}$, and utilize a linear decay scheduler to decay LR to $1\times 10^{-5}$. The overall training process consists of 300,000 steps.
@@ D.1 Structure-aware Generation and Predictive Tasks
Inverse folding (structure-conditioned sequence generation). We report the amino acid recovery (AAR) self-consistency TMscore on the Cameo dataset in Table 13. Our results show that the 650M-parameter DPLM2 variant, with bit-based modeling, outperforms the 3B-parameter baseline. Incorporating geometric architectural designs and representation alignments leads to further improvements. These findings indicate that these design methods contribute positively beyond structure generation.
@@ D.6 Training and Sampling Efficiency
Bit-based vs. hybrid approach. We report the training time of DPLM2 variants with either bit-based modeling or hybrid approach on 16 H100s for 300k training steps in Table 19. The increased training time of flow matching (FM) primarily results from the on-the-fly computation of noisy structure $\mathbf{x}_{t}$ using structure encoder, as we can precompute the discrete structure tokens when FM is not applied. FM remains highly effective when inference efficiency is a priority, as it accelerates the sampling process by 10x by requiring fewer sampling steps.
@@ Appendix E Related Work
Protein language models. Impacted by the success of large language models (LLMs), similar practice has been extended to the development of PLMs. For instance, sequence-based protein language models operate by taking each amino acid sequence as a sentence. ESM-1b (Rives et al., 2019) utilizes self-supervised masked language modeling on 250 million protein sequences spanning evolutionary diversity, later leading to the development of ESM-2 (Lin et al., 2023) that further extends the scaling laws. DPLM (Wang et al., 2024b) adopts the bi-directional receptive field, and an additional discrete diffusion pre-training process which allows DPLM to exhibit a strong generative capability. ProtTrans (Elnaggar et al., 2021), ProteinBERT (Brandes et al., 2022), PRoBERTa (Nambiar et al., 2020), ProtAlbert (Behjati et al., 2022), TAPE (Rao et al., 2019), ProteinLM (Xiao et al., 2021), and CARP (Yang et al., 2022) involve several other representative masked language modeling (MLM) paradigm. These sequence-based PLMs perform competitively with classic methods that rely on multiple sequence alignments, indicating that PLMs have captured some of the evolutionary information from sequences alone. In particular, these protein language models achieve powerful generalization on various downstream tasks involving the secondary and tertiary structures. Recent findings further showcase their capabilities in predicting protein functions (Meier et al., 2021), structure folding (Lin et al., 2023), and de novo designs (Verkuil et al., 2022).
@@
Multimodal protein langauge models. Proteins are defined by amino acid sequences and three-dimensional structures, where the sequence dictates the structure, and the structure, in turn, determines the protein’s function. Due to this intrinsic relationship, there has been increasing interest in developing multimodal PLMs that integrate both sequence and structural information. Many recent research has explored structure-enhanced PLMs, building upon the success of sequence-based PLMs. The key foundation of this direction is the adoption of VQ-VAE (Van Den Oord et al., 2017b), which encodes the 3D structures into discrete tokens that represent the geometric conformation of each protein residue. For instance, SaProt (Su et al., 2023) constructs a structure-aware vocabulary that integrates both 3D and amino acid tokens. ProSST (Li et al., 2024) models the correlations between amino acids and structure tokens through a disentangled attention mechanism. ProstT5 (Heinzinger et al., 2023) extends the sequence-based ProT5 (Elnaggar et al., 2021) by training it to predict structure and sequence tokens conditionally. ESM-3 (Hayes et al., 2024) treats sequence and structures as separate tracks and applies an all-to-all masked language modeling approach. DPLM-2 (Wang et al., 2024a) further advances the joint distribution of sequence and structures through a discrete diffusion mechanism, enabling it to effectively denoise the masked sequence or structure tokens. This method allows DPLM-2 to learn meaningful representations that enhance predictive tasks, while achieving strong generative capabilities, such as conditional sequence generation and co-generation of sequence and structures.
@@
Protein structure generation. Another growing line of work involves structure generation, which aim to design novel protein structures. In this direction, diffusion-based approaches have demonstrated remarkable success. For example, RFDiffusion (Watson et al., 2023) excels at guiding protein structure designs to meet specific functional constraints like enzymes. Chroma (Ingraham et al., 2023) enables the generation of proteins and protein complexes and allows flexible property-based prompting. This approach efficiently generates large structures by transforming collapsed polymers into protein backbones and sequences. ProteinSGM (Lee et al., 2022b) facilitates novel 3D protein design using 2D maps of distances and angles as conditions. Multiflow (Campbell et al., 2024b) achieves structure-sequence co-generation by developing multimodal flow matching that allows the hybrid modeling of continuous structural data and discrete sequences. Despite the promising success in structure generation, these approaches often struggle in sequence-conditioned structure prediction, primarily due to the lack of large-scale training on sequence databases.

### v3

@@ Abstract
Multimodal protein language models (PLMs) integrate sequence and token-based structural information, serving as a powerful foundation for protein modeling, generation, and design. However, the reliance on tokenizing 3D structures into discrete tokens causes substantial loss of fidelity about fine-grained structural details and correlations. In this paper, we systematically elucidate the design space of multimodal PLMs to overcome their limitations. We identify tokenization loss and inaccurate structure token predictions by the PLMs as major bottlenecks. To address these, our proposed design space covers improved generative modeling, structure-aware architectures and representation learning, and data exploration. Our advancements approach finer-grained supervision, demonstrating that token-based multimodal PLMs can achieve robust structural modeling. The effective design methods dramatically improve the structure generation diversity, and notably, folding abilities of our 650M model by reducing the RMSD from 5.52 to 2.36 on PDB testset, even outperforming 3B baselines and on par with the specialized folding models. Project page and code: bytedance.github.io/dplm/dplm-2.1.
@@ 2.1 Multimodal generative learning for proteins with language models and structure tokenization
Diffusion protein language model (DPLM, Wang et al., 2024b), in particular, shows excelling performance in both generation and representation learning of protein sequences, and even structures thanks to its recent multimodal extension DPLM-2 (Wang et al., 2024a). The family of DPLMs grounded in absorbing discrete diffusion framework (Austin et al., 2021; Zheng et al., 2023a), which is characterized by a forward and backward Markov process. Let $\texttt{Cat}(\mathbf{x};\mathbf{p})$ be a categorical distribution on protein sequence $\mathbf{y}$ parameterized by a vector $\mathbf{p}$ on $(|\mathcal{V}|-1)$-dimensional probability simplex. The forward process of discrete diffusion defines a Markov process governed by the transition kernel $q(\mathbf{x}^{(t)}|\mathbf{x}^{(t-1)})=\texttt{Cat}\big{(}\mathbf{x}^{(t)};\beta_{t}\mathbf{x}^{(t-1)}+(1-\beta_{t})\mathbf{q}_{\text{noise}}\big{)}$ that gradually perturb the data $\mathbf{x}\sim q(\mathbf{x})$ into a stationary distribution $\mathbf{x}^{(T)}\sim\mathbf{q}_{\text{noise}}$. The learned backward process $p_{\mathbf{\theta}}(\mathbf{x}^{(t-1)}|\mathbf{x}^{(t)})$ reversely denoises the $\mathbf{x}^{(T)}$ towards the data distribution $\mathbf{x}$, which is typically optimized by the variational bound of the log-likelihood (Ho et al., 2020). The learning objective of absorbing diffusion can be simplified into weighted cross-entropiess, resembling masked language modeling at arbitrary noise levels:
@@ 7 Conclusions
Acknowledgements
We thank anonymous reviewers for their valuable feedback. We would like to especially thank Dr. Hang Li for insightful discussions on the project and feedback on the manuscript that help shape this study, as well as Yi Zhou, Jing Yuan and Yilai Li for their valuable comments.
@@ Appendix B Discussions and Limitations
  • Third, our analysis is conducted on relatively small protein language models (up to 3B), and the scalability of these design choices remains uncertain. We chose this setup to prioritize the analysis of our design space. Further study into how these designs scale with larger models could further enhance multimodal protein language modeling, especially given the proven benefits of scaling in DPLM-2.
@@
Ultimately, the future of multimodal protein foundation models should enable direct atom-level data-space structure modeling; however, this remains an unsolved challenge. Token-based multimodal protein LMs provide a scalable, practical, and effective approach to generative protein modeling. Advancing this approach while addressing its limitations—such as structural information loss, lack of physical constraints, and limited multimodal generalization—will be key to unlocking their full potential and provide crucial insights that help us move closer to achieving atomic-resolution protein modeling in a unified multimodal framework.
@@ B.1 Discussion on Alternative Approaches
  • Alternatives: Training DPLM-2 on continuous tokens. However, it’s unclear whether this approach would work for joint modeling of structures and sequence, which is inherently discrete.
@@ C.2 Bit-level Supervision
In the training stage, we follow Wang et al. (2024a) and utilize the pre-trained DPLM (Wang et al., 2024b) as the parameter initialization to inherit the evolutionary representations learned from the massive protein sequence data, which implicitly captures the structural information that benefits structure modeling (Lin et al., 2022). In addition to the parameters inherited from DPLM, we randomly initialize the structure input-output linear layers. The bit-level DPLM-2 is grounded in the absorbing discrete diffusion framework (Austin et al., 2021; Zheng et al., 2023a). To adapt to the absorbing discrete diffusion framework, we also introduce a learnable absorbing embedding for the absorbing state of structure. We employ 2,000 warmup steps until reaching the maximum learning rate $1\times 10^{-4}$, and utilize a linear decay scheduler to decay LR to $1\times 10^{-5}$. The overall training process consists of 300,000 steps.
@@ D.1 Structure-aware Generation and Predictive Tasks
Inverse folding (structure-conditioned sequence generation). We report the amino acid recovery (AAR) self-consistency TMscore on the Cameo dataset in Table 13. Our results show that the 650M-parameter DPLM-2 variant, with bit-based modeling, outperforms the 3B-parameter baseline. Incorporating geometric architectural designs and representation alignments leads to further improvements. These findings indicate that these design methods contribute positively beyond structure generation.
@@ D.6 Training and Sampling Efficiency
Bit-based vs. hybrid approach. We report the training time of DPLM-2 variants with either bit-based modeling or hybrid approach on 16 H100s for 300k training steps in Table 19. The increased training time of flow matching (FM) primarily results from the on-the-fly computation of noisy structure $\mathbf{x}_{t}$ using structure encoder, as we can precompute the discrete structure tokens when FM is not applied. FM remains highly effective when inference efficiency is a priority, as it accelerates the sampling process by 10x by requiring fewer sampling steps.
@@ Appendix E Related Work
Protein language models. Impacted by the success of large language models (LLMs), similar practice has been extended to the development of PLMs. For instance, sequence-based protein language models operate by taking each amino acid sequence as a sentence. ESM-1b (Rives et al., 2019) utilizes self-supervised masked language modeling on 250 million protein sequences spanning evolutionary diversity, later leading to the development of ESM-2 (Lin et al., 2023) that further extends the scaling laws. DPLM (Wang et al., 2024b) adopts the bi-directional receptive field and scalable discrete diffusion pre-training process (Ye et al., 2023b, a) which allows DPLM to exhibit a strong generative capability. ProtTrans (Elnaggar et al., 2021), ProteinBERT (Brandes et al., 2022), PRoBERTa (Nambiar et al., 2020), ProtAlbert (Behjati et al., 2022), TAPE (Rao et al., 2019), ProteinLM (Xiao et al., 2021), and CARP (Yang et al., 2022) involve several other representative masked language modeling (MLM) paradigm. These sequence-based PLMs perform competitively with classic methods that rely on multiple sequence alignments, indicating that PLMs have captured some of the evolutionary information from sequences alone. In particular, these protein language models achieve powerful generalization on various downstream tasks involving the secondary and tertiary structures. Recent findings further showcase their capabilities in predicting protein functions (Meier et al., 2021), structure folding (Lin et al., 2023), and de novo designs (Verkuil et al., 2022).
@@
Multimodal protein language models. Proteins are defined by amino acid sequences and three-dimensional structures, where the sequence dictates the structure, and the structure, in turn, determines the protein’s function. Due to this intrinsic relationship, there has been increasing interest in developing multimodal PLMs that integrate both sequence and structural information. Many recent research has explored structure-enhanced PLMs, building upon the success of sequence-based PLMs (Zheng et al., 2023b). The key foundation of this direction is the adoption of VQ-VAE (Van Den Oord et al., 2017b), which encodes the 3D structures into discrete tokens that represent the geometric conformation of each protein residue. For instance, SaProt (Su et al., 2023) constructs a structure-aware vocabulary that integrates both 3D and amino acid tokens. ProSST (Li et al., 2024) models the correlations between amino acids and structure tokens through a disentangled attention mechanism. ProstT5 (Heinzinger et al., 2023) extends the sequence-based ProT5 (Elnaggar et al., 2021) by training it to predict structure and sequence tokens conditionally. ESM-3 (Hayes et al., 2024) treats sequence and structures as separate tracks and applies an all-to-all masked language modeling approach. DPLM-2 (Wang et al., 2024a) further advances the joint distribution of sequence and structures through a discrete diffusion mechanism, enabling it to effectively denoise the masked sequence or structure tokens. This method allows DPLM-2 to learn meaningful representations that enhance predictive tasks, while achieving strong generative capabilities, such as conditional sequence generation and co-generation of sequence and structures.
@@
Protein structure generation. Another growing line of work involves structure generation, which aim to generate novel protein structures and functional design (Ye et al., 2025). In this direction, diffusion-based approaches have demonstrated remarkable success. For example, RFDiffusion (Watson et al., 2023) excels at guiding protein structure designs to meet specific functional constraints like enzymes. Chroma (Ingraham et al., 2023) enables the generation of proteins and protein complexes and allows flexible property-based prompting. This approach efficiently generates large structures by transforming collapsed polymers into protein backbones and sequences. ProteinSGM (Lee et al., 2022b) facilitates novel 3D protein design using 2D maps of distances and angles as conditions. Multiflow (Campbell et al., 2024b) achieves structure-sequence co-generation by developing multimodal flow matching that allows the hybrid modeling of continuous structural data and discrete sequences. Despite the promising success in structure generation, these approaches often struggle in sequence-conditioned structure prediction, primarily due to the lack of large-scale training on sequence databases. Recent work further includes sidechains and extend to modeling all-atom structures for generating protein complexes and function design (Chen et al., 2025).